OpenStudy (anonymous):

what weight of water is required to fill a 100 gallon aquarium

OpenStudy (ybarrap):

The density of water is about 1,000.00 kg/m³ There are 264.172 cubic meters in a gallon Weight is mass times gravity. So, $$ \text{mass }=100 \text{ gallons}\times \cfrac{264.172~m^3}{1\text{ gallon}}\times\cfrac{1,000.00~kg}{m^3}\\ \text{weight }_n=\text{mass }\times\text{ gravity}=\text{mass }\times 9.81 \cfrac{m}{s^2} $$ This weight is in Newtons (i.e. \(\text{weight}_n\)) To convert to pounds (i.e. \(\text{weight}_p\)): 1 pound force = 4.44822162 newtons So, in pounds, the weight of 100 gallons of water is, $$ \text{weight}_p=\text{weight}_n\cfrac{1\text{ pound force}}{4.44822162 \text{ Newtons}} $$ That's it!